OpenStudy (anonymous):

The sum of three consecutive odd integers is –27. What are the numbers? A. {–7, –9, –11} B. {–10, –9, –8} C. {–5, –9, –13} D. {–11, –13, –15}

ganeshie8 (ganeshie8):

hint : any two adjacent odd numbers differ by 2

ganeshie8 (ganeshie8):

for ex, 3 and 5 are two adjacent odd numbers, they differ by 5-3 = 2

ganeshie8 (ganeshie8):

so, if u say below are the required odd numbers :- \(x-2,~ x,~ x+2\) then, since they add up to -27, we have :- \(x-2 + x + x+2 = -27\)

ganeshie8 (ganeshie8):

solve \(x\)

OpenStudy (anonymous):

x has to be all the same number right ?

ganeshie8 (ganeshie8):

\(x-2 + x + x+2 = -27\) group like terms \(x+x+x -2 + 2 = -27\) \(3x = -27\) \(x = -9\)

ganeshie8 (ganeshie8):

so, the required numbers wud be :- \(x-2, ~x, ~x+2\) \(-9-2, ~-9, ~-9+2\) \(-11, ~-9, ~-7\)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

Yea it kinda makes sence, but where did this come from. x - 2, x, x + 2 ?

ganeshie8 (ganeshie8):

very good question :) notice that they're differing by "2" so they represent consecutive odd numbers

OpenStudy (anonymous):

and can there be questions like this too ? The sum of three consecutive even integers is –27. What are the numbers? (change odd to even)

OpenStudy (anonymous):

If it's consecutive even integers will it turn into like this. x - 1, x, x + 1?

ganeshie8 (ganeshie8):

let me ask u a different question first :)

ganeshie8 (ganeshie8):

consider sum of "3 even integers"

ganeshie8 (ganeshie8):

can the sum be ever odd ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

then, wat u asking is illegal right ?

ganeshie8 (ganeshie8):

The sum of three consecutive even integers is –27. What are the numbers? (change odd to even)

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

sum of 3 even integers can never be odd, so nobody can find the numbers ;P

ganeshie8 (ganeshie8):

change ur question :)

OpenStudy (anonymous):

You can forget about my last question, thank you :)

ganeshie8 (ganeshie8):

np :)